Analysis of Pricing: You manage The Really Annoying Bicycle Messenger Company

which makes deliveries for lawyers, consultants, accountants, and lobbyists in the much of

Washington DC including Georgetown. The company makes deliveries of documents and small

packages at a price of $10.00 each. The average number of deliveries in one month is 31,000.

The owners of the Really Annoying Bicycle Co. would like to increase its sales and profits. They

know that, if price is lowered, they will generate more deliveries. So they run an experiment.

Price is lowered to $9.00 per delivery in May and the number of deliveries increases to 33,000.

b. Is elasticity elastic, inelastic or neither?
1
Expert's answer
2017-05-04T08:30:08-0400
P1 = $10.00 each, Q1 = 31,000, P2 = $9.00, Q2 = 33,000.
b. The price elasticity of demand is: Ed = (33,000 - 31,000)/(9 - 10)*(9 + 10)/(33,000 + 31,000) = -2,000/1*19/64,000 = -0.59, so the demand for the company services is inelastic.